10-15 pls
010 10.0 points A sinusoidal transverse wave travels along a wire of linear density 8.34 g/m. The wave has amplitude 1.2 cm, frequency 132 Hz and wavelength 3.07 m What is the tension of the wire? Answer in units of N 011 (part 1 of 2) 10.0 points A standing wave is formed on a string that is 32 m long, has a mass per unit length 0.00512 kg/m, and is stretched to a tension of 18 N Find the fundamental frequency Answer in units of cycles/s. 012 (part 2 of 2) 10.0 points Find the next frequency that could cause a standing wave pattern Answer in units of cycles/s. 013 10.0 points A student wants to establish a standing wave on a wire 2.5 m long clamped at both ends. The wave speed is 392 m/s. What is the minimum frequency she should apply to set up standing waves? Answer in units of Hz 014 (part 1 of 2) 10.0 points The length of a string is 1314 cm. It is held fixed at each end. The string vibrates in nine sections: i.e., the string has nine antinodes, and the string vibrates at 140 Hz Find the wavelength. Answer in units of m. 015 (part 2 of 2) 10.0 points What is the fundamental frequency? Answer in units of Hz.
Answer #1

10)

Velocity of transverse wave v=\sqrt{\frac{T}{\mu}} , v=f\lambda

\sqrt{\frac{T}{\mu}}=f\lambda

T=f^2\lambda^2\mu=(132)^2(3.07)^2(8.34*10^{-3})=1370\,N

Tension in the wire is T=1370\,N

----------------------------------------------------------

11)

For a standing wave, fundamental frequency is f=\frac{v}{2L}=\frac{1}{2L}\sqrt{\frac{T}{\mu}}

f=\frac{1}{2*32}\sqrt{\frac{18}{0.00512}}=0.926\,cycles/s

12) Next frequency is f_1=2\frac{v}{2L}=2\frac{1}{2L}\sqrt{\frac{T}{\mu}}=2*\frac{1}{2*32}\sqrt{\frac{18}{0.00512}}=1.852\,cycles/s

------------------------------------------

13) Minimum frequency to set up standing wave is fundamental frequency , f=\frac{v}{2L}

f=\frac{392}{2*2.5}=78.4\,Hz

-----------------------------

14) Wavelength of string is \frac{9\lambda}{2} and length of string is L=1314\,cm

\frac{9\lambda}{2}=L

\lambda=\frac{2L}{9}=\frac{2*1314}{9}=292\,cm

15) Frequency of wave with nine sections as shown in figure is f'=9\frac{v}{2L}=140\,Hz

\Rightarrow  \frac{v}{2L}=15.6\,Hz

Fundamental frequency of wave is f=\frac{v}{2L}=15.6\,Hz

------------------------------------------